Could you please help with this problem?
The profit function for the
recklus Hang gliding service is P(x) = -0.2x^2 +fx - m, where f represents the set up fee for a customer's daily excursion and m represents the monthly hanger rental. Also, P represents the monthly profit in dollars of the small business where x is the number of flight excursions facilitated in a month.
A) if $30 is charged for a set up fee, and the monthly hanger rental is $600, write an equation for the profit, P, in terms of x.
b) How much is the profit when 40 flight excursions are sold in a month?
c) How many flight excursions must be sold in order to maximize the profit? Show the work algebraically.
d) What is the maximum profit.

Answer by htmentor(1343)

The profit function for the
recklus Hang gliding service is P(x) = -0.2x^2 +fx - m, where f represents the set up fee for a customer's daily excursion and m represents the monthly hanger rental. Also, P represents the monthly profit in dollars of the small business where x is the number of flight excursions facilitated in a month.
A) if $30 is charged for a set up fee, and the monthly hanger rental is $600, write an equation for the profit, P, in terms of x.
b) How much is the profit when 40 flight excursions are sold in a month?
c) How many flight excursions must be sold in order to maximize the profit? Show the work algebraically.
d) What is the maximum profit.
======================================================
A) P(x) = -0.2x^2 + 30x - 600
B) P(40) = -0.2*40^2 + 30*40 - 600 = -320 + 1200 - 600 = $280
C) To maximize profit, set dP(x)/dx = 0, solve for x:
dP/dx = -0.4x + 30 = 0 -> x = -30/-0.4 = 75
D) Maximum profit is P(75) = -0.2(75)^2 + 30(75) - 600 = $525